# Junyu Cao and Mohsen Bayati

### v1

@@ 3.2.2 Horizontal moves
Algorithm 2 begins with a prior density $\rho_{0}$ on $[-\beta_{\Theta},\beta_{\Theta}]$ that is positive everywhere. For example, $\rho_{0}$ can be chosen as a uniform distribution. Choose some $p\in(1/2,p_{c})$ and let $q=1-p$. When the algorithm queries at the point $\theta_{k}$ to obtain $Z_{k}$, we update the density $\rho_{k}$:
@@
The next query point $\theta_{k+1}$ is determined as the median of $\rho_{k+1}$. The probability density in the correct side would be increased while the probability density in the wrong side would be reduced. Define $\tau^{\rightarrow}(\delta,a)$ as the number of horizontal moves to reach $a$-neighborhood of $\theta^{*}$ with confidence level $1-\delta$ (which will be specified later). That is, parameter $\theta$ moves to the local $a$-neighborhood of $\theta^{*}$ with probability at least $21-\delta/$ 2 when the number of steps exceeds $\tau^{\rightarrow}(\delta/2,a)$. The horizontal move stops when the total number of movement reaches $\tau^{\rightarrow}(\delta/2,\varepsilon)$ since the precision level $\varepsilon$ is achieved.
@@
    If the horizontal move is less than $\tau^{\rightarrow}(\delta/2,a)$ (indicating it may not yet reach the local neighborhood of $\theta^{*}$ within a distance of $a$), the vertical test stops when it reaches $\tau_{1}^{\uparrow}(\theta)$. Then it moves to the next query point $\theta$;
@@
    If the horizontal move is more than $\tau^{\rightarrow}(\delta/2,a)$ (signifying entry into the local neighborhood of $\theta^{*}$ within a distance of $a$):
@@
      If the vertical move reaches $\tau^{\uparrow}(\delta/2,\varepsilon)$ first, the algorithm terminates, having identified a $\theta$ within the $\varepsilon$ range of $\theta^{*}$;
@@
      Alternatively, if the vertical movement reaches $\tau_{1}^{\uparrow}(\theta)$ first, the vertical test stops and then transitions to the next query point $\theta$.
The rationale behind distinguishing between local and non-local scenarios lies in the observation that two distinct situations can result in a significant number of vertical moves: when both compared models are performing well (i.e., $\theta_{1}$ and $\theta_{2}$ are close to $\theta^{*}$), or when both models are performing poorly (i.e., $\theta_{1}$ and $\theta_{2}$ are far from $\theta^{*}$). In the latter scenario, it is imperative that the algorithm does not terminate prematurely, as $\theta$ has not yet approached $\theta^{*}$ sufficiently. Therefore, the stopping criterion $\tau^{\uparrow}(\delta,\varepsilon)$ is only applied when $\theta$ transitions into the local neighborhood of $\theta^{*}$.
Algorithm 2 is described as follows. We initialize the distribution as $\mu_{0}$. Without any historical data, $\mu_{0}$ can be initialized as the uniform distribution over the confidence region. Alternatively, $\mu_{0}$ can be approximated by the empirical distribution of the estimator from the bootstrap method applying to the first phase. At step $k$, the algorithm tests $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$. If $\theta$ is located at the local neighborhood of $\theta^{*}$ ($k\geq\tau^{\rightarrow}(\delta/2,a)$) and the current step label $s$ is larger than $\tau^{\uparrow}(\delta/2,\varepsilon)$, then the algorithm stops and $\theta=\theta_{k}$. Otherwise, the vertical move stops when $s$ is larger than $\tau_{1}^{\uparrow}(\theta_{k})$. Once the test on $\theta_{k}$ stops, we update prior distribution $\rho_{k}$ and select the next $\theta_{k+1}$ as the median of the updated distribution.
Algorithm 2 APB (Alignment through Probabilistic Bisection)
@@ 3.3 Complexity of Horizontal Moves
Let $T^{\uparrow}(\theta)$ denote the number of comparisons that are collected at $\theta$ and $T^{\rightarrow}$ denote the number of horizontal moves; let $T_{k}^{+}$ denote the total number of comparisons that are collected after $k$ horizontal-moves, where “$+$” denotes both horizontal- and vertical-moves. Note that $T^{\uparrow}(\theta)$, $T^{\rightarrow}$, and $T_{k}^{+}$ are all random variables. In what follows, we discuss the complexity of vertical and horizontal moves, and characterize its sample complexity.
We first characterize the complexity of horizontal moves. Fix some $a\in(0,\theta^{*}+\beta_{\Theta})$. Define three regions: $A=[-\beta_{\Theta},\theta^{*}-a]$, $B=(\theta^{*}-a,\theta^{*}]$, $C=(\theta^{*},\beta_{\Theta}]$. Define $T(a)=\inf\{k^{\prime}\geq 0:\theta_{k}\in B\cup C,\forall k\geq k^{\prime}\}$ to be the time required until the sequence of medians never reenters $A$. Let $\nu_{k}$ denote the (random) measure corresponding to the conditional posterior distribution, conditional on lying to the left of $\theta^{*}$, so that for any measurable $D$, $\nu_{k}(D)=\mu_{k}(D\cap[-\beta_{\Theta},\theta^{*}])/\mu_{k}([-\beta_{\Theta},\theta^{*}])$. Next, fix $\Delta\in(0,1/2)$ and define the stopping time $\tau(a)$ to be the first time that the conditional mass in $B$ lies above $1-\Delta$ and the median $\theta_{k}$ lies to the left of $\theta^{*}$, i.e.,
@@
From (3.1), to bound $T(\cdot)$, we characterize the distribution of $\tau(\cdot)$. Algorithm 2 starts with an initial distribution $\mu_{0}$. Without any prior information, $\mu_{0}$ can be set as a uniform distribution; otherwise we can use the prior information to update $\mu_{0}$. We first make the following mild assumption regarding the initial prior distribution.
@@
The uniform distribution satisfies Assumption 3.5 since $\mu_{0}(B)=\mu_{0}((\theta^{*}-a,\theta^{*}])=a$. With additional information, the prior distribution would be more centered around the true parameter $\theta^{*}$. Thus, $\varsigma$ is assumed to be less than or equal to 1. Under this assumption of prior distribution, we analyze the distribution of $\tau(\cdot)$ in Lemma 3.6.
@@
Lemma 3.6 has shown the light tail distribution of $\tau(\cdot)$. For small $\omega$, note that $\omega^{\varsigma}$ decreases in $\varsigma$. That is, when the initial prior distribution is more centered around the true parameter, $\varsigma$ is smaller and thus the right-hand-side tail distribution is lighter.
@@
Theorem 3.7 states that after $\tau^{\rightarrow}(\delta,\varepsilon)$ horizontal moves, $\theta_{k}$ reaches $\varepsilon$-neighborhood of the true parameter with probability at least $1-\delta$. The complexity of this process depends on $\log(1/\delta)$, $\log(1/\varepsilon)$, and $\varsigma(1+\varsigma)$. When the initial prior distribution is more centered around the true parameter $\theta^{*}$, fewer horizontal moves are needed.
Define the local region of $\theta^{*}$ as $[\theta^{*}-a,\theta^{*}+a]$. We divide the horizontal moves to two phases: Phase I includes all steps until the sequence of medians never reenters $[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,\beta_{\Theta}]$; Phase II includes all steps after Phase I. Similar to the definition of $T(a)$, we define $T^{\prime}(a)$ as the stopping time that the medians never reenters $A^{\prime}=[\theta^{*}+a,\beta_{\Theta}]$. Let $\psi(a)$ be the time required until the sequence of medians never reenters $A\cup A^{\prime}$, i.e., $\psi(a)=\max\{T(a),T^{\prime}(a)\}$. Proposition 3.8 provides an upper bound for the expectation of moves outside of $[\theta^{*}-a,\theta^{*}+a]$.
@@
$\mathbb{E}[\psi(a)]\leq\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}.$
@@ 3.4 Complexity of Vertical Moves
The vertical move has two stopping criteria, depending on whether $\theta_{k}$ has moved to the local neighborhood of $\theta^{*}$. If $\theta_{k}$ is still outside of the local area, the algorithm would finish the power-one test; otherwise, the algorithm may terminate earlier before stopping the power-one test.
Lemma 3.9 first characterizes the expected number of queries needed for testing $\theta$.
@@
Lemma 3.9 implies that the expected number of vertical moves increases in the order of $O((2\tilde{p}(\theta)-1)^{-2})$ when $\tilde{p}(\theta)$ approaches 1/2. When $\tilde{p}(\theta)$ is bounded away from $1/2$, the expected vertical move is also bounded. Lemma 2.6 shows that $\tilde{p}(\theta)$ is strictly larger than 1/2 for any $\theta\neq\theta^{*}$, then according to Heine–Borel theorem, there exists $\delta^{\prime}>0$ such that $\tilde{p}(\theta)>1/2+\delta^{\prime}$ for all $\theta\in[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,-\beta_{\Theta}]$ (see the proof of Lemma 8.9 in Appendix 8.1). For the ease of notation, we define the bound of vertical moves
@@
The selection accuracy bounded away from $1/2$ implies that $\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))$ is bounded. Thus, we can infer the expected number of steps—comprising both horizontal and vertical movements—outside of $[\theta^{*}-a,\theta^{*}+a]$ as follows.
@@
$\mathbb{E}\left[\sum_{k=1}^{\psi(a)}T^{\uparrow}(\theta_{k})\boldsymbol{1}(\|\theta_{k}-\theta^{*}\|\geq a)\right]\leq\left(\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))\,.$
Proposition 3.10 highlights that the sample complexity of Phase I hinges on the choice of $a$ and the span of $\theta$ from the prior distribution. If $\beta_{\Theta}$ is overly expansive, there is a high probability that both selected parameters $\theta_{1}$ and $\theta_{2}$ are distant from the true parameter. Consequently, the accuracy of human selection may diminish considerably, resembling random selection, as illustrated in Example 3.2. Thus, by narrowing the range of $\theta$ based on supplementary information, the sample complexity of Phase I can be notably diminished.
Once Phase I ends, $\theta_{k}$ moves to the $a$-neighborhood of $\theta^{*}$. In scenarios where both choices are close to $\theta^{*}$, discerning the superior parameter becomes more challenging. Depending on whether the selection accuracy converges towards $1/2$ as the detected parameter approaches the true parameter, we classify the problem into two cases: 1) The selection accuracy is bounded away from 1/2 for all $\theta\neq\theta^{*}$; 2) the section accuracy approaches $1/2$. The second case is harder and requires more delicate analysis. We first characterize the sample complexity for case 1) and then analyze case 2).
@@
If $\tilde{p}(\theta)\geq\underline{p}>1/2$ for all $\theta\in\Theta$, then the expected samples collected before algorithm termination is
@@
Next, we address the more challenging scenario in which the selection accuracy tends towards $1/2$ as the tested parameter approaches the true parameter.
@@
As the comparison accuracy approaches $1/2$, the vertical move (power-one test) requires collecting more samples. The rate at which the vertical sample complexity increases depends on the rate at which the utility changes as $\theta$ converges to $\theta^{*}$. To account for this effect, we introduce the following assumption.
@@
The parameter $\kappa$ characterizes the speed of convergence, with a smaller value indicating a faster convergence rate. It is natural to assume $\kappa\leq 1$ for several reasons. Firstly, human decision-making should not be inferior to learning from a random sample; otherwise, collecting random samples would suffice. We demonstrate this by examining the selection accuracy of learning from a single sample. Suppose the sample $(x,y)$ is generated from $y=\theta^{*}x+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. When comparing $\theta+\Delta$ and $\theta-\Delta$ (without loss of generality, assuming $\theta<\theta^{*}$), $\theta+\Delta$ is more preferable than $\theta-\Delta$ if the residual is smaller. That is, the probability of $\theta+\Delta$ being better is
@@
If we use $B$ i.i.d. samples, then the probability of $\theta+\Delta$ being better is $\frac{1}{2}+\sqrt{B}\Omega((\theta^{*}-\theta)|x|)$. The detailed proof is included in Appendix 9.2. In instances where evaluation is conducted solely on a random sample, $\kappa$ is equal to 1. Considering that humans possess additional side information and can make better comparisons, we anticipate $\kappa$ to be, at most, 1.
@@
where $c=\min\{\lambda_{\Delta}/(8\gamma\ln(2)),1/(6a)\}$.
Under Assumption 3.4.1, $\tilde{p}(\theta)$ converges to 1/2 with the speed no faster than the linear convergence. Therefore, when $\theta$ moves to the $a$-neighborhood of $\theta^{*}$, the distance between $\theta$ and $\theta^{*}$ can be bounded by $((\tilde{p}(\theta)-1/2)/c)^{1/\kappa}$. In another word, the closeness between $\tilde{p}(\theta)$ and 1/2 implies the closeness between $\theta$ and $\theta^{*}$. Thus, we can define $\tau^{\uparrow}(\delta,\varepsilon)$ as the threshold of the vertical test such that if the vertical move $T^{\uparrow}(\theta)$ has reached $\tau^{\uparrow}(\delta,\varepsilon)$, then $\theta$ is close enough to $\theta^{*}$.
@@
Suppose Assumption 3.4.1 holds. When $T^{\rightarrow}\geq\tau_{1}^{\rightarrow}(a,\delta/2)$ and $T^{\uparrow}(\theta_{T^{\rightarrow}})\geq\tau^{\uparrow}(\delta/2,\varepsilon)$ where
@@
and $\tau_{0}=\{\max\ s:\hbar_{s}/s\geq c\varepsilon^{\kappa}/2\}$ where $c$ is defined in Lemma 3.12, then we have
@@
When Assumption 3.4.1 holds and $\theta$ moves to the neighborhood of $\theta^{*}$, the sample complexity of the vertical move would not exceed $O(\varepsilon^{-2\kappa})$ where $\varepsilon$ is the precision. The vertical sample complexity increases with the value of $\kappa$. The faster the selection accuracy converges to 1/2, the harder the selection problem, and thus the higher the vertical sample complexity. In the extreme case where $\kappa=0$, $\tilde{p}(\theta)$ is bounded away from 1/2 so $\tau^{\uparrow}(\delta,\varepsilon)$ is bounded. This conclusion is consistent with what we have in Theorem 3.11.

### v2

@@ 4.2.2 Horizontal moves
Algorithm 2 begins with a prior density $\rho_{0}$ on $[-\beta_{\Theta},\beta_{\Theta}]$ that is positive everywhere. For example, $\rho_{0}$ can be chosen as a uniform distribution. Choose some $p\in(1/2,p_{c})$, and let $q=1-p$. When the algorithm queries at the point $\theta_{k}$ to obtain $Z_{k}$, we update the density $\rho_{k}$:
@@
The next query point $\theta_{k+1}$ is determined as the median of $\rho_{k+1}$. The probability density on the correct side would be increased while the probability density on the wrong side would be reduced. Define $\tau^{\rightarrow}(\delta,a)$ as the number of horizontal moves to reach $a$-neighborhood of $\theta^{*}$ with confidence level $1-\delta$ (which will be specifically defined in the paper later). That is, parameter $\theta$ moves to the local $a$-neighborhood of $\theta^{*}$ with probability of at least $21-\delta/$ 2 when the number of steps exceeds $\tau^{\rightarrow}(\delta/2,a)$. The horizontal move stops when the total number of moves reaches $\tau^{\rightarrow}(\delta/2,\varepsilon)$ because at this point the precision level $\varepsilon$ has been achieved.
@@
    If the horizontal move is less than $\tau^{\rightarrow}(\delta/2,a)$ (indicating it may not yet reach the local neighborhood of $\theta^{*}$ within a distance of $a$), the vertical test stops when the movement reaches $\tau_{1}^{\uparrow}(\theta)$. Then it moves to the next query point $\theta$.
@@
    If the horizontal move is more than $\tau^{\rightarrow}(\delta/2,a)$, signifying entry into the local neighborhood of $\theta^{*}$ within a distance of $a$, there are two possible options:
@@
      When the vertical move first reaches $\tau^{\uparrow}(\delta/2,\varepsilon)$, the algorithm terminates, having identified a $\theta$ within the $\varepsilon$ range of $\theta^{*}$.
@@
      When the vertical move first reaches $\tau_{1}^{\uparrow}(\theta)$, the vertical test stops and then transitions to the next query point $\theta$.
The rationale behind distinguishing between local and non-local scenarios lies in the observation that two distinct situations can result in a significant number of vertical moves: (1) when both compared models are performing well (i.e., $\theta_{1}$ and $\theta_{2}$ are close to $\theta^{*}$), or (2) when both models are performing poorly (i.e., $\theta_{1}$ and $\theta_{2}$ are far from $\theta^{*}$). In the latter case, that the algorithm not terminate prematurely is crucial because $\theta$ has not yet gotten sufficiently close to $\theta^{*}$. Therefore, the stopping criterion $\tau^{\uparrow}(\delta,\varepsilon)$ is only applied when $\theta$ transitions into the local neighborhood of $\theta^{*}$.
Algorithm 2 is described as follows. We initialize the distribution as $\mu_{0}$. Without any historical data, $\mu_{0}$ can be initialized as the uniform distribution over the confidence region. Alternatively, $\mu_{0}$ can be approximated by the empirical distribution of the estimator using the bootstrap method and applying it to the first phase. At step $k$, the algorithm tests $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$. If $\theta$ is located at the local neighborhood of $\theta^{*}$ ($k\geq\tau^{\rightarrow}(\delta/2,a)$) and the current step label $s$ is larger than $\tau^{\uparrow}(\delta/2,\varepsilon)$, then the algorithm stops and $\theta=\theta_{k}$. Otherwise, the vertical move stops when $s$ is larger than $\tau_{1}^{\uparrow}(\theta_{k})$. Once the test on $\theta_{k}$ stops, we update the prior distribution $\rho_{k}$ and select the next $\theta_{k+1}$ as the median of the updated distribution.
Algorithm 2 MAPB (Model Alignment through Probabilistic Bisection)
@@ 4.3 Complexity of Horizontal Moves
Let $T^{\uparrow}(\theta)$ denote the number of comparisons that are collected at $\theta$ and let $T^{\rightarrow}$ denote the number of horizontal moves; let $T_{k}^{+}$ denote the total number of comparisons that are collected after $k$ horizontal moves, where “$+$” denotes both horizontal and vertical moves. Note that $T^{\uparrow}(\theta)$, $T^{\rightarrow}$, and $T_{k}^{+}$ are all random variables. In the following paragraphs, we discuss the complexity of vertical and horizontal moves, and we characterize its sample complexity.
We first characterize the complexity of horizontal moves. Fix some $a\in(0,\theta^{*}+\beta_{\Theta})$. Define three regions: $A=[-\beta_{\Theta},\theta^{*}-a]$, $B=(\theta^{*}-a,\theta^{*}]$, and $C=(\theta^{*},\beta_{\Theta}]$. Define $T(a)=\inf\{k^{\prime}\geq 0:\theta_{k}\in B\cup C,\forall k\geq k^{\prime}\}$ to be the time required until the sequence of medians never reenters $A$. Let $\nu_{k}$ denote the (random) measure corresponding to the conditional posterior distribution, conditional on lying to the left of $\theta^{*}$, so that for any measurable $D$, $\nu_{k}(D)=\mu_{k}(D\cap[-\beta_{\Theta},\theta^{*}])/\mu_{k}([-\beta_{\Theta},\theta^{*}])$. Next, fix $\Delta\in(0,1/2)$ and define the stopping time $\tau(a)$ to be the first time that the conditional mass in $B$ lies above $1-\Delta$ and the median $\theta_{k}$ lies to the left of $\theta^{*}$; that is,
@@
From (4.1), to bound $T(\cdot)$, we characterize the distribution of $\tau(\cdot)$. Algorithm 2 starts with an initial distribution $\mu_{0}$. Without any prior information, $\mu_{0}$ can be set as a uniform distribution; otherwise, we can use the prior information to update $\mu_{0}$. First, we make the following mild assumption regarding the initial prior distribution.
@@
The uniform distribution satisfies Assumption 4.5 because $\mu_{0}(B)=\mu_{0}((\theta^{*}-a,\theta^{*}])=a$. With additional information, the prior distribution would be more centered around the true parameter $\theta^{*}$. Thus, $\varsigma$ is assumed to be less than or equal to 1. Under this assumption of prior distribution, we analyze the distribution of $\tau(\cdot)$ in Lemma 4.6.
@@
Lemma 4.6 shows the light tail distribution of $\tau(\cdot)$. For small $\omega$, note that $\omega^{\varsigma}$ decreases in $\varsigma$. That is, when the initial prior distribution is centered more closely around the true parameter, $\varsigma$ is smaller; thus, the right-hand-side tail distribution is lighter.
@@
Theorem 4.7 states that after $\tau^{\rightarrow}(\delta,\varepsilon)$ horizontal moves, $\theta_{k}$ reaches the $\varepsilon$-neighborhood of the true parameter with probability at least $1-\delta$. The complexity of this process depends on $\log(1/\delta)$, $\log(1/\varepsilon)$, and $\varsigma(1+\varsigma)$. When the initial prior distribution is more closely centered around the true parameter $\theta^{*}$, fewer horizontal moves are needed.
Define the local region of $\theta^{*}$ as $[\theta^{*}-a,\theta^{*}+a]$. We divide the horizontal moves into two phases: Phase I includes all steps until the sequence of the medians never reenters $[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,\beta_{\Theta}]$; and Phase II includes all steps after Phase I. Similar to the definition of $T(a)$, we define $T^{\prime}(a)$ as the stopping time, which is when the medians never reenter $A^{\prime}=[\theta^{*}+a,\beta_{\Theta}]$. Let $\psi(a)$ be the time required until the sequence of the medians never reenters $A\cup A^{\prime}$ – that is, $\psi(a)=\max\{T(a),T^{\prime}(a)\}$. Proposition 4.8 provides an upper bound for the expectation of moves outside of $[\theta^{*}-a,\theta^{*}+a]$.
@@
$\mathbb{E}[\psi(a)]\leq\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}.$
@@ 4.4 Complexity of Vertical Moves
The vertical move has two stopping criteria that depend on whether $\theta_{k}$ has moved to the local neighborhood of $\theta^{*}$. If $\theta_{k}$ is still outside of the local area, the algorithm would finish with the power-one test; otherwise, the algorithm may terminate earlier, before stopping the power-one test.
Lemma 4.9 characterizes the expected number of queries needed for testing $\theta$.
@@
Lemma 4.9 implies that the expected number of vertical moves increases in the order of $O((2\tilde{p}(\theta)-1)^{-2})$ when $\tilde{p}(\theta)$ approaches 1/2. When $\tilde{p}(\theta)$ is bounded away from $1/2$, the expected vertical move is also bounded. As a result of Lemma 3.7, $\tilde{p}(\theta)$ is strictly larger than 1/2 for any $\theta\neq\theta^{*}$; thus, according to Heine–Borel theorem, there exists $\delta^{\prime}>0$ such that $\tilde{p}(\theta)>1/2+\delta^{\prime}$ for all $\theta\in[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,-\beta_{\Theta}]$. (See the proof of Lemma 9.9 in Appendix 9.1.) For the ease of notation, we define the bound of vertical moves:
@@
The selection accuracy’s being bounded away from $1/2$ implies that $\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))$ also is bounded. Thus, we can infer the expected number of steps—comprising both horizontal and vertical movements—outside of $[\theta^{*}-a,\theta^{*}+a]$ as follows.
@@
$\mathbb{E}\left[\sum_{k=1}^{\psi(a)}T^{\uparrow}(\theta_{k})\boldsymbol{1}(\|\theta_{k}-\theta^{*}\|\geq a)\right]\leq\left(\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))\,.$
Proposition 4.10 highlights that the sample complexity of Phase I hinges on the choice of $a$ and the span of $\theta$ from the prior distribution. If $\beta_{\Theta}$ is overly expansive, there is a high probability that both selected parameters $\theta_{1}$ and $\theta_{2}$ are distant from the true parameter. Consequently, the accuracy of human selection may diminish considerably – and may even resemble random selection, as illustrated in Example 4.2. Thus, by narrowing the range of $\theta$ based on supplementary information, the sample complexity of Phase I can be notably diminished.
Once Phase I ends, $\theta_{k}$ moves to the $a$-neighborhood of $\theta^{*}$. In scenarios where both options are close to $\theta^{*}$, discerning the superior parameter becomes more challenging. Depending on whether the selection accuracy converges toward $1/2$ as the detected parameter approaches the true parameter, we classify the problem into two cases: 1) The selection accuracy is bounded away from 1/2 for all $\theta\neq\theta^{*}$; and 2) the section accuracy approaches $1/2$. The second case is harder and requires more delicate analysis. We first characterize the sample complexity for Case 1 and then analyze Case 2.
@@
If $\tilde{p}(\theta)\geq\underline{p}>1/2$ for all $\theta\in\Theta$, then the expected number of samples collected before algorithm termination is
@@
Next, we address the more challenging scenario, in which the selection accuracy tends toward $1/2$ as the tested parameter approaches the true parameter.
@@
As the comparison accuracy approaches $1/2$, the vertical move (based on the power-one test) requires the collection of more samples. The rate at which the vertical sample complexity increases depends on the rate at which the utility changes as $\theta$ converges to $\theta^{*}$. To account for this effect, we introduce the following assumption.
@@
The parameter $\kappa$ characterizes the speed of convergence, with a smaller value indicating a faster convergence rate. Assuming that $\kappa\leq 1$ is natural for several reasons. First, human decision making should not be inferior to learning from a random sample; otherwise, collecting random samples would suffice. We demonstrate this by examining the selection accuracy of learning from a single sample. Suppose the sample $(x,y)$ is generated from $y=\theta^{*}x+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. When comparing $\theta+\Delta$ and $\theta-\Delta$ (without loss of generality, assuming $\theta<\theta^{*}$), $\theta+\Delta$ is preferable to $\theta-\Delta$ if the residual is smaller. That is, the probability that $\theta+\Delta$ is better is
@@
In instances where evaluation is conducted solely on a random sample, $\kappa$ is equal to 1. Considering that humans possess or have access to additional side information and thus can make better comparisons, we anticipate that $\kappa$ is, at most, 1.
@@
where $c=\min\{\lambda_{\Delta}/(8\gamma\ln(2)),1/(6a^{\kappa})\}$.
Under Assumption 4.4.1, $\tilde{p}(\theta)$ converges to 1/2 at a speed no faster than the linear convergence. Therefore, when $\theta$ moves to the $a$-neighborhood of $\theta^{*}$, the distance between $\theta$ and $\theta^{*}$ can be bounded by $((\tilde{p}(\theta)-1/2)/c)^{1/\kappa}$. In other words, the closeness between $\tilde{p}(\theta)$ and 1/2 implies the closeness between $\theta$ and $\theta^{*}$. Thus, we can define $\tau^{\uparrow}(\delta,\varepsilon)$ as the threshold of the vertical test, such that if the vertical move $T^{\uparrow}(\theta)$ has reached $\tau^{\uparrow}(\delta,\varepsilon)$, then $\theta$ is close enough to $\theta^{*}$.
@@
Suppose that Assumption 4.4.1 holds. When $T^{\rightarrow}\geq\tau^{\rightarrow}(\delta/2,a)$ and $T^{\uparrow}(\theta_{T^{\rightarrow}})\geq\tau^{\uparrow}(\delta/2,\varepsilon)$, where
@@
and $\tau_{0}=\max\{s:\hbar_{s}/s\geq c\varepsilon^{\kappa}/2\}$, where $c$ is defined in Lemma 4.12, then we have
@@
When Assumption 4.4.1 holds and $\theta$ moves to the neighborhood of $\theta^{*}$, the sample complexity of the vertical move would not exceed $O(\varepsilon^{-2\kappa})$, where $\varepsilon$ is the precision. The vertical sample complexity increases as the value of $\kappa$ increases. The faster the selection accuracy converges to 1/2, the harder the selection problem, and thus the higher the vertical sample complexity. In the extreme case, where $\kappa=0$, $\tilde{p}(\theta)$ is bounded away from 1/2 so $\tau^{\uparrow}(\delta,\varepsilon)$ is bounded. This conclusion is consistent with what we have in Theorem 4.11.